a. A 20 V dc voltage is measured by analog and digital multi-meters. The analog instrument is on 

its 25 V range, and its specified accuracy is ± 2%. The digital meter has a 3½ digital display 

and an accuracy of ± (0.6+1). Determine measurement accuracy in each case.

b. A digital frequency meter has a time base from 1 MHz clock generator frequency divided by 

decade counters. Determine the measured frequency when a 1.512 kHz sine wave is applied 

and the time base uses (a) six decade counters and (b) four decade counters.


1
Expert's answer
2021-05-31T06:20:19-0400

Part a

Analog instrument:

Voltage error "= ±2 \% *25 V = ±0.5 V"

error "=± \frac{0.5 V}{20 V}*100=±2.5\%"


Digital instrument for 20 V displayed on a 3 ½ digit display

1 Digit = 0.1 V

Voltage error = ± (0.6% of reading + 1 Digit)

= ± (1.2 V + 0.1 V)

= ± 0.22 V

error "=± \frac{0.22 V}{20 V}*100=±1.1\%"


Part b

(a) Using six decade counters:

Time base frequency (f1) = "1MHz\/ (10*10*10*10*10*10) = 1Hz"

Time base period (T1) = "1\/f1 = 1\/1Hz = 1 second"

Input frequency period (Ti) = "1\/fi = 1\/1.512kHz"

Cycles counted = "T1\/Ti = T1*fi = 1 second * 1.512kHz = 1512 cycles"

measured frequency on the display = 1.512 kHz.


(b) Using four decade counters:

Time base frequency (f2) = "1MHz\/ (10*10*10*10) = 100Hz"

Time base period (T2) = "1\/f2 = 1\/100Hz = 10ms"

Input frequency period (Ti) "= 1\/fi = 1\/1.512kHz"

Cycles counted ="T2\/Ti = T2*fi = 10ms * 1.512kHz = 15 cycles <lowest integer>"

Measured frequency on the display = 001.5kHz.